### v2

# Typical infinite mixing automorphisms are rank-one

Typical infinite mixing automorphisms are rank-one
Valery V. Ryzhikov
Bashtanov proved that generic mixing automorphisms of probability space with respect to the Alpern-Tikhonov metric had rank one. Using Sidon’s constructions, we show that generic infinite mixing automorphisms also are rank-one.
1 Introduction
Alpern and Tikhonov [1],[7] introduced a complete metric on the space of mixing actions. Bashtanov [2] proved that the class of rank-one actions is generic in the space ${\bf Mix}$ of invertible mixing transformations with invariant probability measure. We will establish the same for the space ${\bf Mix_{\infty}}=\{T:\,T^{n}\to_{w}0\}$ of transformations of $T$ with sigma-finite invariant measure. We prove that the conjugacy class of the Sidon transformation of rank one (see [5], [6]) is dense in ${\bf Mix_{\infty}}$. Since the class of rank-one transformations is a $G_{\delta}$-set, taking into account the density of the latter, we obtain the genericity of this class. A consequence of this and the result of [3] is the typicalness of the minimality of the centralizer (the tipical transformation $T$ commutes only with its powers).
The group ${\bf Aut_{\infty}}$ of all atomorphisms of the space $(X,\cal B,\mu)$ isomorphic to the line ${\mathbf{R}}$ with the standard Lebesgue measure is equipped with the Halmos metric $\rho$:
$\rho(S,T)=\sum_{i}2^{-i}\left(\mu(SA_{i}\Delta TA_{i})+\mu(S^{-1}A_{i}\Delta T^{-1}A_{i})\right),$
where $\{A_{i}\}$ is some fixed family dense in the family of all measurable sets of finite measure. The space $({\bf Aut_{\infty}},\rho)$ is a complete separable metric space. It is easy to show that in $({\bf Aut_{\infty}},\rho)$ the so-called mixing transformations form a set of the first category (countable intersection of nowhere dense sets). If we strengthen the Halmos metric to the Alpern-Tikhonov metric, we obtain a complete metric space ${\bf Mix}_{\infty}$. Which invariants (i.e. properties preserved under conjugation in $\bf Aut_{\infty}$) are typical? A property is called typical if a dense $G_{\delta}$-set of transformations has this property.
Spaces $\bf{\bf Mix}$, ${\bf Mix}_{\infty}$. Following [7], we define the metric $d_{w}$ on ${\bf Aut}(\mu)$:
$d_{w}(S,T)=\sum_{i,j=1}^{\infty}2^{-i-j}\left|\mu(SA_{i}\cap A_{j})-\mu(TA_{i}\cap A_{j})\right|.$
On the set of all mixing automorphisms Mix is defined the metric $r$:
$r(S,T)=\rho(S,T)+\sup_{n>0}d_{w}(S^{n},T^{n}).$
In [7] it is proved that the metric space $({\bf Mix},r)$ is complete and separable.
Consider the sigma-finite measure $\mu_{\infty}$ as an invariant and, without changing the formulas, introduce the same metric $r$ on the space of mixing infinite automorphisms:
${\bf Mix}_{\infty}=\{T\in{\bf Aut}(\mu_{\infty})\,:\,T^{n}\to_{w}0\},$
where $\{A_{i}\}$ is a fixed sequence of sets, dense in the family of all sets of finite measure.
Fundamental in $({\bf Mix_{\infty}},r)$ a sequence $T_{k}$ converges to some $T$ by virtue of the completeness of the Halmos metric in the space ${\bf Aut}(\mu_{\infty})$. It is easy to show that this $T_{k}\to T$ converges to $T$ in the metric $r$ as well.
From the known methods, it can be shown that
for generic transformations in ${\bf Mix}_{\infty}$ that spectra of their symmetric powers are simple.
It follows from the results of [6] that
a typical transformation in ${\bf Mix}_{\infty}$ is not isomorphic to its inverse.
2 Density of the conjugacy classes of the Sidon automorphisms
Let us recall necessary definitions.
Construction of rank one. Fix a natural number $h_{1}$, a sequence $r_{j}\to\infty$ (parameter $r_{j}$ is the number of columns into which the tower of stage $j$ is cut) and a sequence of integer vectors (parameters of spacers)
$\bar{s}_{j}=(s_{j}(1),s_{j}(2),\dots,s_{j}(r_{j}-1),s_{j}(r_{j})),\ s_{(}i)\geq 0.$
At step $j=1$ a set of disjoint half-intervals $E_{1},SE_{1},S^{2}E_{1},\dots,S^{h_{1}-1}E_{1}$ of the same measure is given.
Let at step $j$ a system of non-intersecting half-intervals
$E_{j},SE_{j},S^{2}E_{j},\dots,S^{h_{j}-1}E_{j},$
be defined, and on the half-intervals $E_{j},SE_{j},\dots,S^{h_{j}-2}E_{j}$ the transformation $S$ is the parallel translation. Such a set of half-intervals is called the tower of step $j$, their union is denoted by $X_{j}$ and is also called a tower.
We represent $E_{j}$ as a disjoint union of $r_{j}$ half-intervals
$E_{j}^{1},E_{j}^{2}E_{j}^{3},\dots E_{j}^{r_{j}}$
of the same measure (length). For each $i=1,2,\dots,r_{j}$, we consider the so-called column
$E_{j}^{i},SE_{j}^{i},S^{2}E_{j}^{i},\dots,S^{h_{j}-1}E_{j}^{i}.$
We denote the union of these half-intervals by $X_{i,j}$, which we will also call a column.
To each column $X_{i,j}$ we add $s_{j}(i)$ disjoint half-intervals of the same measure as $E_{j}^{i}$, obtaining a set
$E_{j}^{i},SE_{j}^{i},S^{2}E_{j}^{i},\dots,S^{h_{j}-1}E_{j}^{i},S^{h_{j}}E_{j}^{i},S^{h_{j}+1}E_{j}^{i},\dots,S^{h_{j}+s_{j}(i)-1}E_{j}^{i}$
(all these sets are disjoint). Denoting $E_{j+1}=E^{1}_{j}$, for $i<r_{j}$ we set
$S^{h_{j}+s_{j}(i)}E_{j}^{i}=E_{j}^{i+1}.$
The set of superstructured columns is from now on considered as a tower of stage $j+1$, consisting of half-intervals
$E_{j+1},SE_{j+1},S^{2}E_{j+1},\dots,S^{h_{j+1}-1}E_{j+1},$
where
$h_{j+1}=h_{j}r_{j}+\sum_{i=1}^{r_{j}}s_{j}(i).$
The partial definition of the transformation $S$ at stage $j$ is preserved at all subsequent stages. As a result, on the space $X=\cup_{j}X_{j}$ there is defined an invertible transformation $S:X\to X$ that preserves the standard Lebesgue measure on $X$.
Sidon constructions. Let a rank one construction $S$ have the following property: the intersection $X_{j}\cap S^{m}X_{j}$ for $h_{j}<m\leq h_{j+1}$ can be contained only in one of the columns $X_{i,j}$ of the tower $X_{j}$. Such a construction is called Sidon. The measure of its space $X$ is infinite. If $r_{j}\to\infty$, then Sidon $S$ is mixing.
Theorem 1. Let $T$ be a mixing transformation, $S$ a Sidon mixing transformation of rank one. For every $\varepsilon>0$ there exists a conjugation $R$ such that $r(R^{-1}SR,T)<\varepsilon$.
Lemma 2. Let $T$ be a mixing transformation, $A$ a set of finite measure. The phase space for every natural $H$ can be represented as
$X=\bigsqcup_{h>H}\bigsqcup_{i=0}^{h}T^{i}B_{h},\ \ \ T^{h}B_{h}\cap A=\phi.$
The proof of the lemma is simple, let it be an exercise.
The proof of the theorem (in brief).
Let us consider a part of the Kakutani tower:
$U_{N}=\bigsqcup_{H<h<N}\bigsqcup_{i=0}^{h}T^{i}B_{h},$
which almost absorbed the sets $A_{1},\dots,A_{k}$, appearing in the definition of the metric $r$. For this, we choose a large $N$.
We take a very thin tower $X_{j}$ (so very large $j$), appearing in the construction of the Sidon transformation $S$, and wind on $U_{N}$ the tower $\tilde{X}_{j}$ corresponding to the conjugation $\tilde{S}=R^{-1}SR$ in a way that the following conditions are satisfied:
(1) the conjugate transformation $\tilde{S}$ coincides with $T$ on $U_{N}$ minus the upper floors of $T^{h}B_{h}$ and minus a set of arbitrarily small measure;
(2) after the moving floor of the wound tower leaves the top of $U_{N}$ for the $k$-th time, it must travel outside $U_{N}$ for a time $s_{k}>10s_{k-1}$ ($s_{1}>10N$). This Sidon winding will ensure that the intersection measure is small:
$\mu(\tilde{S}^{m}\tilde{U}_{N}\cap U_{N})\ <\ N\mu(B_{j}),\ \ m<h_{j}.$
For $m\geq h_{j}$, the inequality
$\mu(\tilde{S}^{m}\tilde{U}_{N}\cap\tilde{U}_{N})<\mu(\tilde{U}_{N})/r_{j}$
is satisfied by the definition of the Sidon construction.
If $A=A_{1}\cup\dots\cup A_{k}\subset\tilde{U}_{N}$, then
$\sum_{i}^{k}2^{-i}\left(\mu(SA_{i}\Delta TA_{i})+\mu(S^{-1}A_{i}\Delta T^{-1}A_{i})\right)=0$
and
$\sum_{i,j=1}^{k}2^{-i-j}\left|\mu(SA_{i}\cap A_{j})-\mu(TA_{i}\cap A_{j})\right|=0.$
In reality, for a large $N$ and a suitable Sidon winding, instead of equalities we get inequalities with a very small right-hand side.
So for $\varepsilon>0$ there are $j$, $N$ and $R$ such that at the intersection $\tilde{X}_{j}\cap U_{N}$ the transformations $\tilde{S}$ and $T$ almost coincide, hence, $\tilde{S}$ and $T$ are close in the Halmos metric, but their powers are uniformly close with respect to the $d_{w}$ metric. Thus we obtain $r(R^{-1}SR,T)<\varepsilon$ for a suitable $R$.
The set of rank-one transformations is a $G_{\delta}$ set (use the proof from [2]). We have found that the class of mixing rank one transfomations is dense in $\bf Mix_{\infty}$. Thus, it is typical.
In [4] we used a similar trick, but the windings were random Ornstein and the spacers for flows and automorphisms of rank one were random Ornstein as well.
In the case of infinite measure, a pleasant opportunity arises to use large Sidon spacers, which gives us a simpler and constructive method. For other results on Sidon transformations, see in [5].
3 Типичность ранга один в ${\bf Mix_{\infty}}$
Альперн и Тихонов [1],[7] ввели полную метрику на пространстве перемешивающих действий. Баштанов [2] доказал, что класс действий ранга один типичен в пространстве ${\bf Mix}$братимых перемешивающих преобразований с инвариантной вероятностной мерой. Мы установим аналогичную теорему для пространства ${\bf Mix_{\infty}}=\{T:\,T^{n}\to 0\}$реобразований $T$игма-конечной инвариантной мерой. Докажем, что класс сопряженности сидоновского преобразования ранга один (см. [5], [6]) плотен в ${\bf Mix_{\infty}}$. Так как класс преобразований ранга один является $G_{\delta}$-множеством, с учетом плотности последнего получаем типичность этого класса. Следствием этой типичности с учетом результата [3] является типичность минимальности централизатора (преобразование $T$оммутирует только со своими степенями).
Группа ${\bf Aut_{\infty}}$сех атоморфизмов пространства $(X,\cal B,\mu)$, изоморфного прямой ${\mathbf{R}}$о стандартной мерой Лебега оснащается метрикой Халмоша $\rho$:
$\rho(S,T)=\sum_{i}2^{-i}\left(\mu(SA_{i}\Delta TA_{i})+\mu(S^{-1}A_{i}\Delta T^{-1}A_{i})\right),$
где $\{A_{i}\}$ – некоторое фиксированное семейство, плотное в семействе всех измеримых множеств конечной меры. Пространство $({\bf Aut_{\infty}},\rho)$вляется полным сепарабельным пространством. Несложно показать, что в $({\bf Aut_{\infty}},\rho)$ак называемые перемешивающие преобразования образуют множство первой категории (счетное пересечение нигде не плотных множеств). Если же усилить метрику Халмоша до метрики Альперна-Тихонова, получим полное метрическое пространство ${\bf Mix}_{\infty}$). Возникает вопрос: какие инварианты, т.е. свойства, сохранящиеся при сопряжении в $\bf Aut_{\infty}$) являются типичными? Свойство называется типичным, если плотное $G_{\delta}$-множество преобразований обладают этим свойством.
Например, с учетом известных методов можно показать, что в ${\bf Mix}_{\infty}$ипична однократность спектра всех симметрических степеней преобрзования. Из результатов [6] легко вытекает, что типичное в ${\bf Mix}_{\infty}$реобразование не изоморфно своему обратному.
Пространства $\bf{\bf Mix}$, ${\bf Mix}_{\infty}$. Cледуя [7], определим на ${\bf Aut}(\mu)$етрику $d_{w}$:
$d_{w}(S,T)=\sum_{i,j}2^{-i-j}\left|\mu(SA_{i}\cap A_{j})-\mu(TA_{i}\cap A_{j})\right|.$
На множестве всех перемешиващих автоморфизмов Mix задаем метрику $r$:
$r(S,T)=\rho(S,T)+\sup_{n>0}d_{w}(S^{n},T^{n}).$
В [7] доказано, что метрическое пространство $({\bf Mix},r)$вляется полным и сепарабельным.
Рассмотрим в качестве инвариантной сигма-конечную меру $\mu_{\infty}$, не меняя формул, введем такую же метрику $r$а пространстве перемешивающих бесконечных автоморфизмов:
${\bf Mix}_{\infty}=\{T\in{\bf Aut}(\mu_{\infty}),\,T^{n}\to_{w}0\},$
где $\{A_{i}\}$ – фиксированная последовательность множеств, плотная в семействе всех множеств конечной меры.
Фундаментальная в $({\bf Mix_{\infty}},r)$оследовательность $T_{k}$ходится к некоторому $T$илу полноты метрики Халмоша в пространстве ${\bf Aut}(\mu_{\infty})$. Несложно показать, что в этом случае $T_{k}\to T$етрике $\rho$.
Плотность класса сопряженности сидоновского автоморфизма.
Конструкция ранга один.
Фиксируем натуральное число $h_{1}$, последовательность $r_{j}\to\infty$ ( параметр $r_{j}$ – число колонн, на которые разрезается башня этапа $j$) и последовательность целочисленных векторов (параметров надстроек)
$\bar{s}_{j}=(s_{j}(1),s_{j}(2),\dots,s_{j}(r_{j}-1),s_{j}(r_{j})).$
На шаге $j=1$о определению задан набор полуинтервалов $E_{1},SE_{1},S^{2}E_{1},\dots,S^{h_{1}-1}E_{1}$. Пусть на шаге $j$пределена система непересекающихся полуинтервалов
$E_{j},SE_{j},S^{2}E_{j},\dots,S^{h_{j}-1}E_{j},$
причем на полуинтервалах $E_{j},SE_{j},\dots,S^{h_{j}-2}E_{j}$ребразование $S$вляется параллельным переносом. Такой набор полуинтервалов называется башней этапа $j$, их объединение обозначается через $X_{j}$оже называется башней.
Представим $E_{j}$ак дизъюнктное объединение $r_{j}$олуинтервалов
$E_{j}^{1},E_{j}^{2}E_{j}^{3},\dots E_{j}^{r_{j}}$
одинаковой меры (длины). Для каждого $i=1,2,\dots,r_{j}$ассмотрим так называемую колонну
$E_{j}^{i},SE_{j}^{i},S^{2}E_{j}^{i},\dots,S^{h_{j}-1}E_{j}^{i}.$
Объдинение этих полуинтервалов обозначим через $X_{i,j}$, его тоже будем называть колонной.
К каждой колонне $X_{i,j}$обавим $s_{j}(i)$епересекающихся полуинтервалов той же меры, что у $E_{j}^{i}$, получая набор
$E_{j}^{i},SE_{j}^{i},S^{2}E_{j}^{i},\dots,S^{h_{j}-1}E_{j}^{i},S^{h_{j}}E_{j}^{i},S^{h_{j}+1}E_{j}^{i},\dots,S^{h_{j}+s_{j}(i)-1}E_{j}^{i}$
(все эти множества не пересекаются). Обозначив $E_{j+1}=E^{1}_{j}$, для $i<r_{j}$оложим
$S^{h_{j}+s_{j}(i)}E_{j}^{i}=E_{j}^{i+1}.$
Набор надстроеных колонн с этого момента рассматривается как башня этапа $j+1$, состоящая из полуинтервалов
$E_{j+1},SE_{j+1},S^{2}E_{j+1},\dots,S^{h_{j+1}-1}E_{j+1},$
где
$h_{j+1}=h_{j}r_{j}+\sum_{i=1}^{r_{j}}s_{j}(i).$
Частичное определение преобразования $S$а этапе $j$охраняется на всех следующих этапах. В итоге на пространстве $X=\cup_{j}X_{j}$пределено обратимое преобразование $S:X\to X$, сохраняющее стандартную меру Лебега на $X$.
Пусть конструкция $S$анга один обладает следующим свойством: пересечение $X_{j}\cap S^{m}X_{j}$ри $h_{j}<m\leq h_{j+1}$ожет содержаться только в одной из колонн $X_{i,j}$ашни $X_{j}$. Такая конструкция называется сидоновской. Мера пространства $X$том случае бесконечна.
Теорема 1. Пусть $T$ – перемешивающее преобразование, $S$ – сидоновское преобразование ранга один. Для всякого $\varepsilon>0$айдется такое сопряжение $R$, что $r(R^{-1}SR,T)<\varepsilon$.
Лемма 2. Пусть $T$ – перемешивающее преобразование, $A$ – множество конечной меры. Фазовое пространство для всякого натурального $H$ожно представить в виде
$X=\bigsqcup_{h>H}\bigsqcup_{i=0}^{h}T^{i}B_{h},\ \ \ T^{h}B_{h}\cap A=\phi.$
Доказательство леммы несложное, оставляем его в качестве упражнения.
Доказательство теоремы кратко. Рассмотрим часть башни Какутани:
$U_{N}=\bigsqcup_{H<h<N}\bigsqcup_{i=0}^{h}T^{i}B_{h},$
которая почти поглотила множества $A_{1},\dots,A_{k}$, фигурирующие в определении метрики $r$. Для этого выбираем большое $N$.
Возьмем очень тонкую башню $X_{j}$, фигурирующую в построении сидоновского преобразования $S$, намотаем на $U_{N}$оответствующую сопряжению $\tilde{S}=R^{-1}SR$ашню $\tilde{X}_{j}$аким образом, чтобы выполнялись условия:
(1) сопряженное преобразование $\tilde{S}$овпадало с $T$а $U_{N}$инус верхние этажи $T^{h}B_{h}$инус множество сколь угодно малой меры;
(2) после того как движущийся этаж намотанной башни покидает в $k$-тый раз верх $U_{N}$, он должен путешествовать вне $U_{N}$ремя $s_{k}>10s_{k-1}$ ($s_{1}>10N$). Этот способ можно назавать сидоновской намоткой, которая обеспечит нам малость меры пересечения:
$\mu(\tilde{S}^{m}\tilde{U}_{N}\cap U_{N})\ <\ N\mu(B_{j}),\ \ m<h_{j}.$
При $m\geq h_{j}$ыполняется неравенство
$\mu(\tilde{S}^{m}\tilde{U}_{N}\cap\tilde{U}_{N})<\mu(\tilde{U}_{N})/r_{j}$
в силу определения сидоновской конструции.
Если бы выполнялось $A=A_{1}\cup\dots\cup A_{k}\subset\tilde{U}_{N}$, то
$\sum_{i}^{k}2^{-i}\left(\mu(SA_{i}\Delta TA_{i})+\mu(S^{-1}A_{i}\Delta T^{-1}A_{i})\right)=0$
и
$\sum_{i,j=1}^{k}2^{-i-j}\left|\mu(SA_{i}\cap A_{j})-\mu(TA_{i}\cap A_{j})\right|=0.$
В действительности при выборе большого $N$одходящей обмотке вместо равенств получим неравенства с очень малой правой частью.
Итак для $\varepsilon>0$сегда найдутся $j$, $N$ $R$акие, что на пересечении $\tilde{X}_{j}\cap U_{N}$реобразования $\tilde{S}$ $T$очти совпадают, значит близки в метрике Халмоша, но их степени равномерно близки относительно метрики $d_{w}$. Тем самым мы получим $r(R^{-1}SR,T)<\varepsilon$ля подходящего $R$.
Множество преобразований ранга один является $G_{\delta}$ножеством (используем без особых изменений доказательство из [2]). Мы показали его плотность. Таким образом, типичность ранга один доказана.
Замечание. В [4] мы использовали подобный прием, но обмотки были случайными орнстейновскими и надстройки для потоков и автоморфизов ранга один также были случайными орнстейновскими. В случае бесконечной меры возникает приятная возможность использовать сидоновкие надстройки, что дает простой и конструктивный метод. Другие результаты о сидоновских преобразованиях см. в [5].
Список литературы
  • [1] Alpern St., Conjecture: In general a mixing transformation is not two-fold mixing, Ann. Probab., 13:1 (1985), 310-313
  • [2] А.И. Баштанов, Типичное перемешивание имеет ранг 1, Матем. заметки, 93:2 (2013), 163-171; A.I. Bashtanov, Generic Mixing Transformations Are Rank 1, Math. Notes, 93:2 (2013), 209-216
  • [3] В.В. Рыжиков, Ж.-П. Тувено, О централизаторе бесконечного перемешивающего преобразования ранга один, Функц. анализ и его прил., 49:3 (2015), 88-91; V.V. Ryzhikov, J.-P. Thouvenot, On the Centralizer of an Infinite Mixing Rank-One Transformation, Funct. Anal. Appl., 49:3 (2015), 230-233
  • [4] В.В. Рыжиков, О сохраняющих меру преобразованиях ранга один, Тр. ММО, 81, № 2, МЦНМО, М., 2020, 281-318; V.V. Ryzhikov, Measure-preserving rank one transformations, Trans. Moscow Math. Soc., 81:2 (2020), 229-259
  • [5] В.В. Рыжиков, Полиномиальная жесткость и спектр сидоновских автоморфизмов. Матем. сб., 215:7 (2024), 138-152; V.V. Ryzhikov, Polynomial rigidity and spectrum of Sidon automorphisms, Mat. Sb., 215:7 (2024), 138-152
  • [6] V.V. Ryzhikov, Asymmetric mixing Poisson suspensions, Arxiv 2406.14390
  • [7] С.В. Тихонов, Полная метрика на множестве перемешивающих преобразований, Матем. сб., 198:4 (2007), 135-158; S.V. Tikhonov, A complete metric in the set of mixing transformations, Sb. Math., 198:4 (2007), 575-596